what is the slope of a line that is perpendicular to the graph of: y=-1/2x+4?
Found 2 solutions by nerdybill, bucky:
Answer by nerdybill(7384)

since the slope of:
y=-1/2x+4
is -1/2
.
a line perpendicular to this must have a slope that is the "negative reciprocal":
(-1/2)m = -1
(1/2)m = 1
m = 2
.
answer is 2

Answer by bucky(2189)

This problem can be done in your head if you know the method.
.
If you know the slope of a given line, the line perpendicular to it will have a slope that is the negative inverse.
.
In this problem the known slope is negative 1/2. So the line that is perpendicular to it will have a slope that is positive and has a slope of the inverse of 1/2 which is 2/1 or just 2.
.
So the answer to this problem is a slope of +2.
.
Hope this helps you to understand how to find the slope of lines perpendicular to a line of known slope. Just find the negative inverse.